OpenStudy (anonymous):

10. Every month, a girl gets allowance. Assume last year she had no money, and kept it up to now. Then she spends 1/2 of her money on clothes, then 1/3 of the remaining money on games, and then 1/4 of the remaining money on toys. After she bought all of that, she had $7777 left. Assuming she only gets money by allowance, how much money does she earn every month?

OpenStudy (anonymous):

please help!!!

OpenStudy (anonymous):

Let x be her monthly allowance assume that a year has 12 months so she gets 12x if 1 represents all her money from 12x she spends 1/2 so another 1/2 of the original remain = 0.5 she then spends 1/3 so 2/3 of the half remain (1/2)*(2/3) = 2/6 = 1/3 than spends another 1/4 of the that remainder so 3/4 of that reminder remains. so (1/3) * (3/4) = 3/12 = 1/4 remains from the 1 so if 1/4 = $7777 1 = 4 * 7777 = $31108 so now if she got 12 months of allowance 12x = $31108 x = ?

OpenStudy (anonymous):

2222